0:30 WFS powered up. Some noise on OCAM top left and bootm right quadrants when powering up but rebinning resolved.

OCSO

Star RA05h11+14, catalog magnitude R=7.53

1:54 Pointing and collimation binocular (in preparation for science). Seeing looks to be 0.5-0.8" on the guiders to start. Some thin cirrus

1:55 Reauthorizing LUCI1 only for OSCO

1:58 OSCO preset sent for Star RA05h11+14 HD33607, seeing is 0.6" on the guider.
  • OCAM magnitude 7.9
  • Bin1 900Hz OCAM gain 500, 81.2 photons/subapp,
  • gains tt3.22, 0.62, 0.73 pre optical gain, optical gain and NCPA on 1.31, 0.25, 0.29 with an optical gain ~0.42
  • Loop opened after just a few short seconds of NCPA applied.
  • Reoptimizing: 2.93, 0.67, 0.60, then optical gain and as soon as NCPA is applied we get the bite at 9O'clock
  • TN while starting NCPA: TN20220313_021107,021117 just before bite takes over, rotator at 154.
  • luci1.20220313.0061-0080 NO NCPA
  • 1st POSITION, NCPA OFF
    => Command line arguements processed
    Set 0 Tracknum 20220313_021430 LuciName unknown
    Set 1 Tracknum 20220313_021443 LuciName unknown
    Set 2 Tracknum 20220313_021455 LuciName unknown
    -> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220313_021430.log

    2nd POSITION, NCPA OFF
    => Command line arguements processed
    Set 0 Tracknum 20220313_021716 LuciName unknown
    Set 1 Tracknum 20220313_021728 LuciName unknown
    Set 2 Tracknum 20220313_021740 LuciName unknown
    -> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220313_021716.log

AZ Science

HOPS 299

ESM Target Catalog magnitude AOref = 11.23, elevation

7:23 Preset sent for HOPS 299. AO rotator out of allowed range. Steve rotating us 360 and resending preset

2:26 Preset resent after rotator put in happy place.
  • OCAM magnitude 12.7, 7.6 photons/subapp bin1 freq400, some thin cirrus
  • Gains 7.43, 6.33, 0.08
  • TN20220313_023621
  • We have signitficat moon illumination on the LUCI1 GCS. About 41 degrees moon distance with moon at 75% and cirrus.

HOPS 102

ESM target with catalog magnitude 15.9
  • Optimize gain just safe skipps open when ramping TT, 1
  • The moon background is huge and this AOref is within a nebula. . Not much correction here. TT1.1, H01 0.55 was best i could do based on modal plot.
  • TN 033437
  • TN 034253 seeing 0.84"
  • WfsArbitrator _178759|INF| 969|2022-03-13 03:51:37.948762| MAIN > offsetXY(): starting offset of 0.093, -17.797 mm
    WfsArbitrator_178759|INF| 970|2022-03-13 03:51:37.949561| MAIN > offsetXYZ(): Focus correction (absolute): 0.000 mm
    WfsArbitrator_178759|INF| 971|2022-03-13 03:52:18.007829| MAIN > Stage movement timeout after 40 seconds. Retry 1 of 3
    WfsArbitrator_178759|INF| 972|2022-03-13 03:53:03.546346| MAIN > Stage movement timeout after 40 seconds. Retry 1 of 3
    WfsArbitrator_178759|INF| 973|2022-03-13 03:53:46.648987| MAIN > Stage movement timeout after 40 seconds. Retry 2 of 3
    WfsArbitrator_178759|ERR| 974|2022-03-13 03:54:29.751520| MAIN > RETRY: Stages position not reached (-5001) TIMEOUT_ERROR

  • Sending a (0,0) offset. (Alex accidentally sent relative)
  • 977|2022-03-13 03:56:43.427841| MAIN > offsetXYZ(): Focus correction (absolute): 0.000 mm
    WfsArbitrator_178759|INF| 978|2022-03-13 03:57:23.494559| MAIN > Stage movement timeout after 40 seconds. Retry 1 of 3
    WfsArbitrator_178759|INF| 979|2022-03-13 03:58:06.631102| MAIN > Stage movement timeout after 40 seconds. Retry 2 of 3
    WfsArbitrator_178759|ERR| 980|2022-03-13 03:58:49.737214| MAIN > RETRY: Stages position not reached (-5001) TIMEOUT_ERROR

  • Y stage requested position

    CopleyCtrl _178565 |ERR| 529|2022-03-13 03:58:09.667159| MAIN > Request: -85.442104 Limits: 0.000000 -85.000000
    CopleyCtrl_178565 |ERR| 530|2022-03-13 03:58:09.667245| MAIN > Variable handler error: [AOException] Error in MoveTo (): movement out of range (code -5002) Value is out of range:

  • (0,0) absolute offset recovered the AO just fine. Size of original offset was that failed (0, -20.37).
  • 4:30 Same thing with a Y stage failure, offset commanded (0,-20.37)

    9|INF| 1045|2022-03-13 04:29:32.338597| MAIN > Stage movement timeout after 40 seconds. Retry 1 of 3
    WfsArbitrator_178759|INF| 1046|2022-03-13 04:30:15.435165| MAIN > Stage movement timeout after 40 seconds. Retry 2 of 3
    WfsArbitrator_178759|ERR| 1047|2022-03-13 04:30:58.533472| MAIN > RETRY: Stages position not reached (-5001) TIMEOUT_ERROR

    CopleyCtrl_178565 |ERR| 688|2022-03-13 04:28:52.279135| MAIN > Error: movement out of range
    CopleyCtrl_178565 |ERR| 689|2022-03-13 04:28:52.279215| MAIN > Request: -85.511813 Limits: 0.000000 -85.000000
    CopleyCtrl_178565 |ERR| 690|2022-03-13 04:28:52.279308| MAIN > Variable handler error: [AOException] Error in MoveTo(): movement out of range (code -5002) Value is out of range
    CopleyCtrl_178565 |ERR| 691|2022-03-13 04:29:35.374237| MAIN > Error: movement out of range
    CopleyCtrl_178565 |ERR| 692|2022-03-13 04:29:35.374333| MAIN > Request: -85.511813 Limits: 0.000000 -85.000000
    CopleyCtrl_178565 |ERR| 693|2022-03-13 04:29:35.374431| MAIN > Variable handler error: [AOException] Error in MoveTo(): movement out of range (code -5002) Value is out of range
    CopleyCtrl_178565 |ERR| 694|2022-03-13 04:30:18.471074| MAIN > Error: movement out of range
    CopleyCtrl_178565 |ERR| 695|2022-03-13 04:30:18.471164| MAIN > Request: -85.511813 Limits: 0.000000 -85.000000
    CopleyCtrl_178565 |ERR| 696|2022-03-13 04:30:18.471259| MAIN > Variable handler error: [AOException] Error in MoveTo(): movement out of range (code -5002) Value is out of range
    CopleyCtrl_178565 |INF| 697|2022-03-13 04:32:01.749317| MAIN > Moving to: -84.104 mm
Sending absolute (0,0) resolved

Ilocater

4:50 Going to Polaris
  • Some cirrus, Polaris -1.46, OCAM magnitude 1.2. Seeing 1.02" on the guider at 32 degrees elevation.
  • 20220313_050455. Optical gain 0.317, no NCPA
  • Ocam slope rms/ave pupil intensity, incoinglight, blah blah are no longer updating. The loop clains to be closed but target walking off.
  • Rebinning fixed the OAM population issue but real time pupils did not return, even after a new HO dark
  • resending preset AO was the ulitmate fix
  • Applygin Z5=8, a little Z10 would be nice but loop seems a bit unstable
  • bin1 freq, mod3, ocam gain 7
  • ADC test successful, proved that the images look better with the ADC on.
Slewed to Peocyon for final ADD test. Verified that NCPA on does induce open fram every 10 seconds on Ilocater camera.

OSCO

6:09 Slewing to OSCO target #11 BD+70 622 Catalog mag R7.48
  • Elevated noise in teh upper right and lower left quadrands but does not appear to impact closing.
  • OCAM magnitude 8.7
  • Rotator 179, el 53
  • Seeing 1.1,
  • TN 061819 Just after NCPA On. There was already an edge threatening to furl before NCPA on. Right after this was taken the edge curled and was a bite at 10'Oclock and opened.
  • Running with NO NCPA luci1.20
06:26:21 Failure on set disturb, when optimizing gain
  • adsc_stop/start. (verified lbti wfs off prior to the restarting of the adsec) Flat set with 535 modes.
  • 6:36 preset resent
  • Camera lens not moving with out manual intervention during close loop. Doug tried to manually kick it and restarting the pupil process but no luck. Soptting preset and restarting the WFS arb
  • 6:44 Reseting preset , 1.1". the loop closed and edge did not curl. Image looksed good. luci1.20220313.0120.
  • 6:53 Loop opened with curled edge, rotatoer at 192.

Repoint test

6:55 Sending preset to same osco target to test repoint. Repoint radio button selected in AOS Commands.
  • Repoint was selected, LUCI script seems to have deselected it?
  • Baysides at: -73.4 -47.5
  • There is some uncertainly if repoint is selected
  • Center star is moving bayside stages with preset. May need to send presets with idl
Doug is sending and IDL preset to AOStar 392 with GS#2, PA 90
  • Tests show the Repoints click on the AOS GUI but needs to be part of the preset. The repoint select should be visible on all AOS gui's open.
  • 1) 7:05 IRC command for runAO with repoint command
    • Baysides at before repoint -76.28 -50.26, then teh X bayside stage at -77.95 and -50.262 after repoint.
  • 2) 7:13 AcquireRef though AOS command GUI rather than IRC command
    • -Baysides at 76.27, -50.264 before center star, and at -76.992 and -51.441 after center star.
  • So repoint function is completely broken
  • WfsArbitrator_117459|INF|       539|2022-03-13 07:15:47.992035|             MAIN > source_acq.acquire(): starting source acquisition WfsArbitrator_117459|INF|       540|2022-03-13 07:15:47.993921|             MAIN > Exception: thGetVar wfsarb.L.CLICK_ACQUIRE error(-3006): Variable not found error - Assuming automatic acquisition WfsArbitrator_117459|INF|       541|2022-03-13 07:15:52.505395|             MAIN > Star found at x=147.5, y=84.7, flux=9089.6 WfsArbitrator_117459|INF|       542|2022-03-13 07:15:52.507200|             MAIN > source_acq.acquire(): star found at X=590.10,Y=338.79 WfsArbitrator_117459|INF|       543|2022-03-13 07:15:52.507315|             MAIN > source_acq.acquire(): current error X=0.798 mm, Y=1.306 mm WfsArbitrator_117459|INF|       544|2022-03-13 07:16:03.637915|             MAIN > Star found at x=122.1, y=123.7, flux=6750.7 WfsArbitrator_117459|INF|       545|2022-03-13 07:16:03.639691|             MAIN > source_acq.acquire(): star found at X=488.41,Y=494.74 WfsArbitrator_117459|INF|       546|2022-03-13 07:16:03.639754|             MAIN > source_acq.acquire(): current error X=0.003 mm, Y=0.088 mm WfsArbitrator_117459|INF|       547|2022-03-13 07:16:03.639789|             MAIN > source_acq.acquire(): target reached. WfsArbitrator_117459|INF|       548|2022-03-13 07:16:03.640744|             MAIN > offsetXYZ(): Focus correction (absolute): 0.000 mm WfsArbitrator_117459|INF|       549|2022-03-13 07:16:03.737662|             MAIN > WARNING: TTOffset LUT disabled due to issues with TT mirror offsets WfsArbitrator_117459|DEB|       550|2022-03-13 07:16:03.737953|             MAIN > arbFunc wrapper done.

Bin3 Faint

RA08hrs Star 15, PA 120

7:25 Slew to bin3 faint test star. Catalog magnitude,
  • OCAM Estimated magnitude 16.1
  • 12-20 Average pupil intensity
  • Bin 3, Freq 345, OCAM gain 600. Gain 5.11, 1.47, 0.1 then apply optical gain. GOpt 0.079
  • TN 20220313_073555
  • LUCI images are a bit elongated 0.09"

-- JenniferPower - 13 Mar 2022
  • modal_1.jpg:
    modal_1.jpg
  • Gains 5.51, 1.87, 0.07 without optical gain. This was verified with the modal plots and with slope rms (~0.74 without optical gain). Still a touch elongated
  • No optical gain: TN075403
  • With Optical gain: TN075814
  • With NCPA: 080028, Z5 ~5 and Z6 ~-8. Looks better, rounder and less elongated.
8:06 Resending the bin3 target as 16.6 to force bin4
  • Closed with bin4, freq 452,
  • bin4 optimize gain plots actually had minimums.
  • Running with optical gain 0.1 TN 081331
  • The modal plot suggests that bin 4 might be doing a bit more.... The bin4 images do look rounder but seeing has gone up so we can't eaily do a comparison
  • 35-45 average pupil intensity.
  • Modal plot of target pushed from bin3 to bin4:
    Bin4_push_modal.png

Bright End Bin3 target

8:19 Slewing to star 10 in RA10-11hr, catalog mag is 13.92
  • Target should be fainter...but OCAM showed ti to be 13.5 during center pupils which puts it in bin2
  • Moving on
8:23 Slewing to star 7 in RA12-13hr Catalog 14.2
  • OCAM magnitude 15.9
  • Bin3 Freq 451: Optimize gain with tt 5.41, 1.27, 0.10, 15-20 average pupil intensity
  • Optimize gain, no optical gain or NCPA TN#083121
  • Optimized and Optical gain with GOpt 0.11 TN 083401
  • NCPA on , Z5 is 34, Z6 is 38, TN 083515
  • PSF looks a little ugly with some trefoil even with NCPA, lui1.20220313.0132
08:38 Slew to star6 in RA14-15hr, starRmag 13.5
  • bin2 shows upper right and lower left quadrant noise
  • OCAM magnitude is 13.9 so moving on
  • Switching binnings before we move on to clean up noisy quadrants. OCAM went not live, switch back resolved. And noise is cleaned up
08:44 Slewing to star 7 in RA14-15hr, Catalog mag R=14.2
  • OCAM magnitude 15.5, average pupil intensity ~15. Bin3, frequency 750.
  • Optimize gain, optimized 4.10, 1.34, 0.10, no Gopt, no NCPA:TN085357
  • Optical Gain on (Gopt 0.075) TN085517
  • Taking some data without NCPA to start.
  • TN 091217
  • 9:13 Turning NCPA on, X5 55.6 , Z6 -33.1 0.9"
  • With NCPA on: TN091644, luci1.20220313.0144-48 (including sky offsets) Images do look a lot better with NCPA on. Can see some airy rings in bin 3, but some instability in the cube 0.9"
  • with NCPA 0.84" TN 093028

Close binary test

9:35 Slewing to close binary target
  • rotator out of range, steve is moving us 360
9:37 HIP65069 , AO ref catalog magnitude 6.8
  • companion 0.12" within 1.5 mags of primary
  • bin1 900Hz ocam gain 401. Rotator 348, estimated OCAM mag 8.4
  • Optimize gain 3.58, 0.88, 0.71, Optical gain on - GOpt 0.385
  • NCPA on Z5~-19, Z6 ~31 TN#094619
9:59, large offset while safeskipping. Safe skip appeared at 100% while in pause AO. When we returned to source we did not resume AO. Was able to resume AO from AOS
  • Looop instabilities, ~8% safeskip TN 100020, TN 1000149
10:03 Reoptimizing with optical gain of 0.345. No NCPA to rule that out as a source of instability
  • TN 100550 to start
  • I have not noted any loop instabilities stince turning NCPA off so that was likely the cause and not the close binary
10:26 Turning back on NCPA
  • The pendingSkipFrame does increment every 10 seconds like it does for LBTI when NCPA is on
  • TN103102. Didnt see loop instabilities this time but have to handover to AZ
10:30 Safing the WFS, warming, and shutting down.
Topic revision: r7 - 13 Mar 2022, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback